OpenStudy (anonymous):

HELP PLEASE complete the square to write the equation in standard form x^2+y^2-8x+10y+15=0

OpenStudy (anonymous):

x² + y² - 8x + 10y +15=0 (x² -8x +4²) +( y² + 10y + 5²) = -15 + 16 + 25 ( x - 4)² + ( y + 5)² = 26